OpenStudy (anonymous):

can somebody find the root of x(2)-x-6=0

myininaya (myininaya):

(x-3)(x+2)=0 x=3 x=-2

OpenStudy (anonymous):

do you know what the quadratic equation of 10x(2)=x=2 using a graph is

OpenStudy (anonymous):

10x(2)+x=2 i mean

myininaya (myininaya):

its a parabola

OpenStudy (anonymous):

what is the root of[9x-8=x^{2}

OpenStudy (anonymous):

what is the root of x^{2}-7x-4=10